Solve the pfaffian D.E. y(x+4)(y+z)dx - x(y+3z)dy+2xydz =0


1
Expert's answer
2021-03-29T14:12:21-0400

"y(x+4)(y+z)dx-x(y+3z)dy+2xydz=0\\
X=(P,Q,R)\text{ --- vector field}.\\
X\cdot \text{curl } X=0\\
X=(y(x+4)(y+z),-x(y+3z),2xy)\\
\text{curl }X=det
\begin{bmatrix}
\overline{i} & \overline{j} & \overline{k}\\
\frac{\partial}{\partial x} & \frac{\partial}{\partial y} & \frac{\partial}{\partial z}\\
P & Q &R
\end{bmatrix}=\\
=5x\overline{i}+(2y+yx)\overline{j}+(-9y-7z-2xy-xz)\overline{k}\\
\text{We can check that }X\cdot \text{curl } X=0.\\
\text{Let }x=const.\text{ Then:}\\
-x(y+3z)dy+2xydz=0|:(-x)\\
(y+3z)dy-2ydz=0|:dy\\
y+3z-2yz^{\prime}=0|:2y\neq 0\\
z^{\prime}=\frac{y+3z}{2y}\\
z^{\prime}=\frac{1}{2}+\frac{3}{2}\cdot\frac{z}{y}\\
t(y)=\frac{z}{y};\ z=t\cdot y;\ z^{\prime}=t^{\prime}\cdot y+t.\\
t^{\prime}\cdot y+t=\frac{1}{2}+\frac{3}{2}t\\
t^{\prime}\cdot y-\frac{1}{2}\cdot t-\frac{1}{2}=0\\
\frac{dt}{dy}y-\frac{1}{2}t-\frac{1}{2}=0\\
ydt-\frac{1}{2}tdy-\frac{1}{2}dy=0\\
ydt+(-\frac{1}{2}t-\frac{1}{2})dy=0|:y(-\frac{1}{2}t-\frac{1}{2})\neq 0\\
\frac{dt}{-1/2t-1/2}+\frac{dy}{y}=0\\
-2\frac{dt}{t-1}+\frac{dy}{y}=0|\int\\
-2\ln|t-1|+\ln|y|=\ln C\\
\ln\frac{|y|}{(t-1)^2}=\ln C\\
\frac{y}{(t-1)^2}=C\\
U(x,y,z)=\frac{y}{(z/y-1)^2}=C.\\
\exist \mu:\ \mu Q=\frac{\partial Q}{\partial y}.\\
\frac{\partial Q}{\partial y}=\frac{-z-y}{y(z/y-1)^7}\\
\mu=\frac{-z-y}{y(z/y-1)^7}\cdot \frac{1}{x(y+3z)}=\frac{z+y}{xy(y+3z)(z/y-1)^7}.\\
k=\mu P-\frac{\partial U}{\partial x}.\\
\frac{\partial U}{\partial x}=0\\
k=\mu P\\
k=\frac{(y+z)^2(x+4)}{x(y+3z)(z/y-1)^7}\\
\frac{\partial U}{\partial x}+k=0\\
\int kdx=C\\
U(x,y,z)=\frac{(x+4)(y+z)^2}{x(y+3x)(z/y-1)^7}=C."